OpenStudy (anonymous):

Write the first derivative of the given function. f(x)=[ ln(x) ]^2 / x^2

OpenStudy (anonymous):

Use this: d(f(x)/(g(x)))/ dx = (g(x)f'(x)-f(x)g'(x))/g(x)^2, where f'(x) = df/ dx denotes the derivative of f with respect to x.

OpenStudy (anonymous):

I did that much I got \[x ^{2}\times2lnx-[\ln(x)]^{2}\times2x \div x(x ^{4)}\]

OpenStudy (anonymous):

i just don't see how to simply to the answer

OpenStudy (anonymous):

yes - i slipped up - ill check it out on paper

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

use quotient rule x^2 * 2lnx *(1/x) - 2x (lnx)^2 ------------------------- x^4 = 2x lnx - 2x(ln x )^2 -------------------- x^4 = 2xlnx(1 - lnx) ----------- x^4

OpenStudy (anonymous):

the answer is \[2(lnx)(1+lnx) \div x ^{3}\]

OpenStudy (anonymous):

and where did 1/x go?

OpenStudy (anonymous):

oh yes - i didnt cancel out the x and x^4 - that gives the answer you just quoted the 1/x cancelled out with x^2 in the first line

OpenStudy (anonymous):

ok and how did you get 1-lnx?

OpenStudy (anonymous):

1-lnx is correct thats the quotinet rule 1 + lnx is incorrect

OpenStudy (anonymous):

quotient rule if y = u/v then dy/dx =[ vdu/dx - udv/dx ] / v^2